OpenStudy (anonymous):

What is the 33rd term of this arithmetic sequence? 12, 7, 2, -3, -8, … A. -158 B. -157 C. -148 D. -147

OpenStudy (anonymous):

NEED HELP

OpenStudy (haseeb96):

here a =12 d = 7-12= -5 while n= 33 we have a formula an= a+(n-1)d a33= 12 + (33-1)(-5) = 12 +(32)(-5) now solve it yourself and get your answer

OpenStudy (haseeb96):

thank you

OpenStudy (anonymous):

thanks